OpenStudy (anonymous):

Roger is 5 feet tall and casts a shadow 3.5 feet long. At the same time, the flagpole outside his school casts a shadow 14 feet long. Write and solve a proportion to find the height of the flagpole.

OpenStudy (anonymous):

h/ 14 = 5/ 3.5 Use cross product: h = ...

OpenStudy (anonymous):

yeah; ratio of similar triangles do what chlorophyll said